# supplementary_materials

### v1

@@ I Introduction
The recent advent of artificial, synthetic platforms have provided a fertile ground for the controlled exploration of a myriad of strongly-correlated electronic phases of matter Andrei et al. (2021); Nuckolls and Yazdani (2024). Remarkably, in a number of synthetic multilayer graphene and transition metal dichalcogenide (TMD) systems, fascinating phenomena – including electron nematicity, exciton condensates, superconductivity, strange metallicity, and quantum anomalous Hall effects – have already been discovered, in some cases in a single system (for a sample of papers, see Refs. Cao et al. (2018a, b); Sharpe et al. (2019); Serlin et al. (2020); Wu et al. (2021); Cao et al. (2021); Oh et al. (2021); Xie et al. (2021); Polshyn et al. (2019); Cao et al. (2020); Jaoui et al. (2022); Chen et al. (2020); Nuckolls et al. (2023); Lu et al. (2024); Han et al. (2024); Tang et al. (2019); Regan et al. (2020); Wang et al. (2020); Li et al. (2021a); Huang et al. (2021); Li et al. (2021b); Ghiotto et al. (2021); Zhao et al. (2023); Cai et al. (2023); Park et al. (2023); Tao et al. (2024); Kang et al. (2024); Liu et al. (2024)). These systems afford a level of tunability and control that is arguably unparalleled, especially when compared to archetypal strong-correlated systems where one is largely restricted to chemical doping or pressure tuning for the more stoichiometrically rigid systems. Synthetic platforms, on the other hand, possess additional knobs, such as an external electric (displacement) field or an intra-layer dependent bias potential, that can be applied, which open a new avenue into the types of phases that can be realized as well as possible phase transitions Song et al. (2024). Compounded with the variety of heterostructure combinations, a rich variety of correlated phases can be realized and studied in a deliberate fashion, which may help to shine light on the outstanding questions in the field of strong-correlated electronic systems.
@@
Motivated to examine synthetic platforms that realize proximate phases and phase transitions, we focus on spatially separated doped bilayer semiconductors with strong electrically tunable interlayer correlations. This metallic setup is conducive to interlayer exciton condensation. Indeed, such a setting is particularly fertile as (i) the exciton condensate phase can be uniform in real-space or have additional translational symmetry breaking (with a finite-momentum ordering order parameter), and (ii) can potentially host a continuous exciton condensation phase transition. This finite-momentum ordering is reminiscent of the finite-$\mathbf{Q}$ spin-density wave phase transition, which has a storied and admittedly complicated abundance of studies (see Refs. Löhneysen et al. (2007); Lee (2018) and references therein). In our study, we focus on type-II band aligned bilayer semiconductors separated by an insulating barrier, that are dual-gated and separately contacted, allowing for the application of an interlayer bias voltage $V_{b}$. Such a system enables conduction electrons and holes to reside natively in their own separated layers, thus forming a $p-n$ junction. Here, type-II indicates a ‘staggered’ alignment of the conduction and valence bands from each layer Lo et al. (2011); Özçelik et al. (2016); Sun et al. (2021); Kistner-Morris et al. (2024) that is favorable to the formation of interlayer excitons. We consider systems having $C_{2z}$ symmetry, and therefore for our calculations, we pick the example case of WS2/WTe2 heterostructure, even though our results hold more generally for a bilayer heterostructure.
We consider charge-imbalanced situations, and find that the metallic state is unstable to a finite-momentum (incommensurate) exciton condensation, as has been noted in previous works Pieri et al. (2007); Varley and Lee (2016). Crucially, this instability can be electrically controlled via a bias potential ($V_{b}$) and carrier density ($n$), as shown in the phase diagram in Fig. 2 employing self-consistent Hartree-Fock calculations. We refer to the exciton condensed phase as the exciton density wave (X-DW) Bi and Fu (2021). We analyze Goldstone mode mediated interactions between fermions in the X-DW phase and show that they can be attractive leading to superconductivity. Our Hartree-Fock calculations are suggestive of a quantum critical line that separates the X-DW from the normal metal with no exciton condensate. Sufficiently close to the quantum critical line, we show that low-energy exciton modes mediate interactions that favor a pair-density wave (PDW) superconductor at the same-ordering momentum as the X-DW, over a zero-momentum pairing. We discuss experimental probes that are sensitive to these phenomena, highlighting in particular, drag transport measurements.
II Microscopic Model of $p-n$ Junction
@@
Figure 1: Left: Schematic of semiconductor -- hBN spacer -- semiconductor sandwich. The densities on the individual layer are controlled independently with an external bias potential $V_{b}$ applied staggerdly on each layer. Right top: Schematic of the unequally occupied conduction and valence bands in the electron-hole plasma state. Right bottom: Schematic of the Fermi surfaces, where the red (blue) is an electron (hole)-like Fermi surface. The exciton density wave order is triggered by the condensation of electron-hole pairs connected by the wave-vector $|\mathbf{Q}|$.
@@
We consider a gate-tunable $p-n$ junction structure, wherein two semiconducting monolayers are separated by an insulating barrier, as shown in Fig. 1. The top and bottom layers are biased externally such that the native bandgap of the individual semiconductors are overcome by an applied displacement field or bias voltage. We will hereafter measure the bias voltage relative to the value required to achieve band touching, and label it $V_{b}$. Moreover, the monolayers are connected to their own electronic reservoir Zheng et al. (1997); Zeng and MacDonald (2020); Tu et al. (2024), which enables the electronic density on each layer to be individually tuned. Specifically, we examine the situation where the top/bottom layers are electron/hole doped. The corresponding kinetic energy of the conduction and valence electrons is,
$H^{0}_{c/v}$ $=\sum_{\mathbf{k}}\xi_{c/v,\mathbf{k}}f^{\dagger}_{c/v,\mathbf{k}}f_{c/v,\mathbf{k}},$ (1)
where the conduction and valence band dispersions are given by $\xi_{c,\mathbf{k}}=\left(\frac{k_{x}^{2}}{2m_{c,x}}+\frac{k_{y}^{2}}{2m_{c,y}}-\frac{V_{b}}{2}-\mu\right)$ and $\xi_{v,\mathbf{k}}=-\left(\frac{k_{x}^{2}}{2m_{v,x}}+\frac{k_{y}^{2}}{2m_{v,y}}-\frac{V_{b}}{2}+\mu\right)$, where $m_{c/v,x/y}$ are the anisotropic effective masses of the electrons in the conduction (top) and valence (bottom) layers, and $\mu$ is the chemical potential that fixes the total carrier density $n$.
@@
Within each layer there is in-principle a two-fold spin degeneracy (or alternatively valley degeneracy as the valley is locked to spin via spin-orbit coupling Manzeli et al. (2017)), which introduces additional possibilities for the spin-correlations of the exciton condensate. We circumvent this complication, by drawing motivation from a recent experimental study Nguyen et al. (2023), where spin-polarized excitons have been reported in Coulomb-coupled monolayers of MoSe2 and WSe2; indeed mean-field studies of charge-compensated ${\mathbf{Q}=0}$ exciton condensates also indicate that spin-polarization is stabilized by an infinitesimal Zeeman field Wu et al. (2015). As such, we focus on an analogous situation, where a spin-polarized exciton condensate forms (for example, spin-$\uparrow$ from the conduction/top layer and spin-$\downarrow$ from the valence/bottom layer). In this sense, the spin-layer labels are ‘locked’. With this understanding, we henceforth drop the spin-labels and keep the layer labels in what follows.
@@
+.$\Big{.}2e^{-qd}n_{c}(\mathbf{q})n_{v}(-\mathbf{q})\Big{]},+ .$ (2)
where $A$ is the area of the system, $V(\mathbf{q})=\frac{e^{2}}{2\epsilon_{0}\epsilon q}\text{tanh}(qd_{g})$ is the screened Coulomb interaction for a dual-gated setup with an inter-gate distance $d_{g}$, and interlayer separation $d$. We have compactly defined the conduction and valence densities, $n_{c/v}(\mathbf{q})=\sum_{\mathbf{k}}f^{\dagger}_{c/v,\mathbf{k}+\mathbf{q}}f_{c/v,\mathbf{k}}$. The system has a combined $U(1)_{c}\times U(1)_{v}$ symmetry, respectively corresponding to the independent conservation of particle number in each layer; alternatively, the sum-of-charges in each layer ($n_{c}+n_{v}$) – with corresponding symmetry $U(1)_{t}$ – and difference-of-charges in each layer ($n_{c}-n_{v}$) – with corresponding symmetry $U(1)_{r}$, are also conserved. Our low-energy continuum model of the $p-n$ junction also implies that continuous translations along the two directions are good symmetries, which become modified to lattice translational symmetries once a lattice potential is introduced. The interaction terms on the first line of Eq. II describe the intra-layer interactions, while the last term denotes the inter-layer interactions. Recent experiments Wang et al. (2019); Ma et al. (2021); Shi et al. (2022); Nguyen et al. (2023) appear to realize inter-layer excitonic instabilties, which suggests that intra-layer interactions play the benign role of renormalizing the dispersion (and effective masses) and the native bandgap of the semiconductors; as such its role in triggering additional (within each layer) electronic instabilities is disregarded.
III Exciton density wave Order
@@
The non-intersecting electron and hole Fermi surfaces in the respective top and bottom layers are susceptible to an exciton-formation instability. In particular, with an unequal number of electron and hole carriers, the respective circular (elliptical) electron(hole) Fermi surfaces are separated by finite-momentum, which leads to the formation of finite-momentum excitons with a preferred center-of-mass momentum, $\mathbf{Q}$, as depicted in Fig. 1. This physical intuition for the directional dependence of the center-of-mass momentum is explicitly confirmed in Appendix C, where the determined center-of-mass momentum direction is extracted from examining the momentum location of the peak of the interlayer susceptibility of the system within a random phase approximation (RPA). We find that the RPA susceptibility is peaked when the exciton center-of-mass momentum is ±$\pm\mathbf{Q}$, where $\mathbf{Q}=\pm(k_{F,v,x}-k_{F,c,x})\hat{x}$ where $k_{F,v/c,x}$ are Fermi momenta in the two layers along the $x$-direction, i.e. the vector connecting the electron-hole Fermi surfaces along the $\hat{x}$-direction.
[…]
@@
Figure 3: Development of collinear exciton density wave for $n=-0.7\times 10^{12}$ cm-2 and $V_{b}=12$ meV within a three-band Hubbard approximation (valence band for ${\mathbf{k}\pm\mathbf{Q}}$ and conduction band for ${\mathbf{k}}$). Left: Electronic Fermi surface Fermi surface in the collinear X-DW phase (yellow denotes the filled Fermi sea). Right: Electronic Bandstructure for $k_{x}=0$ in the presence/absence of the collinear exciton condensate in red/black.
IV Collective Mode mediated superconductivity
The presence of a symmetry-broken exciton condensate phase offers the possibility of examining instabilities triggered by the associated excitonic collective modes. In this section, we will specifically focus on the development of superconducting instabilities mediated by collective modes: (i) gapless Goldstone modes within the broken-symmetry phase, and (ii) the critical mode associated with the phase transition. For analytical tractability of our analysis, we assume that the gate-screened Coulomb interactions become over-screened in the limit of small gate distances $d_{g}\ll 1/q$, leading to a momentum-independent Hubbard interaction.
On general grounds, interlayer exciton fluctuations couple locally to the fermions (in real space $\mathbf{r}$) as,
[…]
IV.1 Goldstone fluctuations mediated nodal superconductivity
@@
In the collinear X-DW condensate, generators of $U(1)_{r}$ and continuous translation along the $\hat{x}$-direction are independently broken, leading to two independent Goldstone modes [^1] We note that in the case of the spiral X-DW, a combination of the generators remains as a symmetry, for example, the $U(1)_{r}$ phase transformation, $\chi=\chi_{+}e^{i(\mathbf{Q}\cdot\mathbf{r})}\rightarrow\chi_{+}e^{i(\mathbf{Q}\cdot\mathbf{r}+\theta)}$, can be compensated by a lattice translation with $\alpha=-\theta$. This compensation entails the breaking of effectively one symmetry generator leading to one Goldstone mode. associated with: a uniform $U(1)_{r}$ phase rotation, $\chi=\chi_{0}\cos(\mathbf{Q}\cdot\mathbf{r})\rightarrow e^{i\theta}\chi_{0}\cos(\mathbf{Q}\cdot\mathbf{r})$, and a uniform translation $\chi=\chi_{0}\cos(\mathbf{Q}\cdot\mathbf{r})\rightarrow\chi_{0}\cos(\mathbf{Q}\cdot\mathbf{r}+\alpha)$. While the coupling of these Goldstone modes to the fermions retains the integrity of the Fermi liquid quasiparticles Watanabe and Vishwanath (2014), we ask here whether they can mediate superconductivity. A similar theoretical study was performed in the context of inter-valley coherent fluctuation-mediated superconductivity in twisted bilayer graphene Po et al. (2018); Kozii et al. (2022) and multilayer graphene Chatterjee et al. (2022); Vituri et al. (2024).
@@
We confine ourselves to the consideration of static fluctuations of the collinear X-DW order parameter: $\chi(\mathbf{r})=\chi_{0}\cos({\mathbf{Q}\cdot\mathbf{r}})+\delta\chi_{-}(\mathbf{r})e^{-i\mathbf{Q}\cdot\mathbf{r}}+\delta\chi_{+}(\mathbf{r})e^{i\mathbf{Q}\cdot\mathbf{r}}$, where the first term is the mean-field (saddle-point) solution. Retaining the important phase fluctuations ($\text{Im}[\delta\chi_{\pm}(\mathbf{r})]\neq 0$) at low energies, we obtain a Gaussian action for these fluctuations following the standard recipe of integrating out high-energy fermions (see Appendix D for details). Diagonalizing the $2\times 2$ fluctuation Hamiltonian, we obtain two gapless Goldstone modes: (i) ‘in-phase’ fluctuation mode ($\delta\chi_{+}=\delta\chi_{-}$, labelled by $\delta\chi_{s}$), which can be interpreted as a superfluid mode due to the breaking of $U(1)_{r}$, and (ii) an ‘out-of-phase’ mode ($\delta\chi_{+}=-\delta\chi_{-}$, labelled by $\delta\chi_{p}$), which can be interpreted as an acoustic phonon due to breaking of the continuous-translation symmetry by the fluctuations. The corresponding Goldstone mode Hamiltonian is,
$H_{G}=\frac{\rho_{G}}{2}\sum_{\mathbf{q}}\left(q_{x}^{2}+\kappa_{G}q_{y}^{2}\right)\delta\chi_{G}(-\mathbf{q})\delta\chi_{G}(\mathbf{q}),$ (4)
@@
where $G=s/p$ labels the two (normal) Goldstone modes, $\rho_{G}$ is the Goldstone mode stiffness and $\kappa_{G}$ is the anisotropy parameter, both of which are numerically calculated. We note that we have disregarded the dynamics of the Goldstone mode here, which when treated provides a more accurate estimates of the critical temperature (within the framework of Eliashberg theory). Since our focus here is on the qualitative behaviors of the superconductivity, we focus solely on the static components; similar studies of fluctuation mediated superconductivity also demonstrate the predictive power of examining the purely static component of the Goldstone action Po et al. (2018); Kozii et al. (2022); Vituri et al. (2024). Employing the defined Goldstone modes ($\delta\chi_{G}$) in terms of the bare fluctuations ($\delta\chi_{\pm}$), we thus obtain the interaction term between the low-energy fermions $f$ (formed from the mean-field band) and the Goldstone modes:
$H_{f-G}=-i\int_{\mathbf{q},\mathbf{k}}\alpha_{G}(\mathbf{k},\mathbf{q})\delta\chi_{G}(-\mathbf{q})f^{\dagger}_{\mathbf{k}+\mathbf{q}}f_{\mathbf{k}}.$ (5)
Finally, we integrate out the Goldstone modes to obtain an effective interaction term between the low-energy fermions:
$\sum_{G}\frac{1}{2A\rho_{G}}\sum_{\mathbf{k},\mathbf{k}^{\prime},\mathbf{q}}f^{\dagger}_{\mathbf{k}+\mathbf{q}}f^{\dagger}_{\mathbf{k}^{\prime}-\mathbf{q}}f_{\mathbf{k}^{\prime}}f_{\mathbf{k}}\frac{\alpha_{G}(\mathbf{k},\mathbf{q})\alpha_{G}(\mathbf{k}^{\prime},-\mathbf{q})}{q_{x}^{2}+\kappa_{G}q_{y}^{2}}.$ (6)
@@
As described in detail in Appendix D, the effective interaction mediated by $\delta\chi_{p/s}$ in the BCS channel, for small momentum transfer is attractive/repulsive, respectively. Therefore we focus our attention on the phonon collective mode $\delta\chi_{p}$ in the following, and decouple the interaction term in the BCS channel and self-consistently solve for the superconducting gap function. We obtain superconducting solutions, i.e., the order parameter $\frac{1}{N_{k}}\left(\sum_{\mathbf{k}}|\langle f_{-,\mathbf{k}}f_{-,-\mathbf{k}}\rangle|^{2}\right)^{1/2}\neq 0$, for sufficiently large $d_{g}$ (equivalently, small $\epsilon$). Here $N_{k}$ is the number of momentum points used in the mesh. Figure 4 depicts the superconducting order parameter (top) and the BCS gap function $\Delta(\mathbf{k})$ (bottom) for a representative parameter point. The gap function has an anisotropic p-wave character, resulting in nodal points in the BdG spectrum. Further, the superconducting state inherits the time-reversal symmetry breaking of the normal X-DW state. While time-reversal symmetry broken superconductors are uncommon in nature, there is recent evidence for this phenomenon in rhombohedrally-stacked tetra-layer graphene Han et al. (2024).
@@
Figure 4: Goldstone mode ($\delta\chi_{p}$)-mediated superconductivity inside the collinear X-DW phase. Top: Superconducting order parameter as a function of $V_{b}$ for $n=-0.7\times 10^{12}$ cm-2. Bottom: BCS gap function as a function of momentum for $V_{b}=12$ meV. Results were obtained using a $59\times 59$ momentum mesh, $\epsilon=10$, momentum space cutoff along the $x$-direction $k_{c,x}=0.2$.
IV.2 Critical fluctuations mediated pair-density wave
@@
Our mean-field theory results in Fig. 2 suggest a continuous phase transition between the metallic electron-hole plasma and the X-DW along a critical line. The region near this line is thus ripe for the development of further electronic instabilities. In the standard lore of metallic quantum critical points, superconductivity and non-Fermi liquids are two of the competing instabilities Löhneysen et al. (2007); Sachdev (2011). Since our goal is not to study this competition in detail Metlitski et al. (2015), we focus here on the possibility of unconventional superconductivity arising from pairing of the fermions on approach from the Fermi liquid. In Appendix F, we consider non-Fermi liquid behavior from the classic one-loop hot-spot model treatment which leads to a dynamical critical exponent $z=3$, and singular corrections of the electronic self-energy, $\Sigma(\omega)\sim i|\omega|^{2/3}$. We emphasize here that our results that follow do not rely on there being a strictly continuous phase transition – it suffices to have a weakly first-order transition with a sufficiently long correlation length, as will be shown.

### v2

# Unconventional superconductivity mediated by exciton density wave fluctuations

@@ Abstract
The recent advent of artificial, synthetic platforms have provided a fertile ground for the controlled exploration of a myriad of strongly-correlated electronic phases of matter. Remarkably, in a number of synthetic multilayer graphene and transition metal dichalcogenide (TMD) systems, fascinating phenomena – including electron nematicity, exciton condensates, superconductivity, strange metallicity, and quantum anomalous Hall effects – have already been discovered, in some cases in a single system (for a sample of papers see Refs. [11, 12, 58, 57, 75, 13, 46, 76, 52, 10, 25, 15, 44, 37, 22, 65, 53, 71, 33, 24, 32, 19, 79, 9, 48, 66, 27, 34]); for overarching reviews on graphene and TMD based moiré materials see Refs. [40, 4, 45] ). These systems afford a level of tunability and control that is arguably unparalleled, especially when compared to archetypal strong-correlated systems where one is largely restricted to chemical doping or pressure tuning for the more stoichiometrically rigid systems. Synthetic platforms, on the other hand, possess additional knobs, such as an external electric (displacement) field or an intra-layer dependent bias potential, that can be applied, which open a new avenue into the types of phases that can be realized as well as possible phase transitions [61]. Compounded with the variety of heterostructure combinations, a rich variety of correlated phases can be realized and studied in a deliberate fashion, which may help to shine light on the outstanding questions in the field of strongly-correlated electronic systems.
@@
Motivated to examine synthetic platforms that realize proximate phases and phase transitions, we focus on spatially separated doped bilayer semiconductors with strong electrically tunable interlayer correlations. This metallic setup is conducive to interlayer exciton condensation. Indeed, such a setting is particularly fertile as (i) the exciton condensate phase can be uniform in real-space or have additional translational symmetry breaking (with a finite-momentum ordering order parameter), and (ii) can potentially host a continuous exciton condensation phase transition. This finite-momentum ordering is reminiscent of the finite-$\mathbf{Q}$ spin-density wave phase transition, which has a storied and admittedly complicated abundance of studies (see Refs. [36, 31] and references therein). In our study, we focus on generic type-II band aligned bilayer semiconductors separated by an insulating barrier (such as a WS2/WTe2 heterostructure), that are dual-gated and separately contacted, allowing for the application of an interlayer bias voltage $V_{b}$. Such a system enables conduction electrons and holes to reside natively in their own separated layers, thus forming a pnp-n junction. Here, type-II indicates a ‘staggered’ alignment of the conduction and valence bands from each layer [35, 47, 63, 29] that is favorable to the formation of interlayer excitons.
@@
Figure 1: Left: Schematic of semiconductor - hBN spacer - semiconductor sandwich. The densities on the individual layer are controlled independently with an external bias potential $V_{b}$ applied staggerdly on each layer. Right top: Schematic of the unequally occupied conduction and valence bands in the electron-hole plasma state. Right bottom: Schematic of the Fermi surfaces, where the red (blue) is an electron (hole)-like Fermi surface. The exciton density wave order is triggered by the condensation of electron-hole pairs connected by the wave-vector $|\mathbf{Q}|$.
@@
Microscopic model of pnp-n junction.— We consider a gate-tunable pnp-n junction structure, wherein two semiconducting monolayers are separated by an insulating barrier, as shown in Fig. 1. The top and bottom layers are biased externally to overcome the native bandgap of the individual semiconductors. We will hereafter measure $V_{b}$ relative to the value required to achieve band touching. Moreover, the monolayers are connected to their own electronic reservoir [80, 77, 67], which enables the electronic density on each layer to be individually tuned. Specifically, we examine the situation where the top/bottom layers are electron/hole doped. The corresponding kinetic energy is, $H^{0}_{c/v}=\sum_{\mathbf{k}}\xi_{c/v,\mathbf{k}}f^{\dagger}_{c/v,\mathbf{k}}f_{c/v,\mathbf{k}}$ , where the conduction and valence band dispersions are $C_{2z}$ symmetric and given by $\xi_{c,\mathbf{k}}=\left(\frac{k_{x}^{2}}{2m_{c,x}}+\frac{k_{y}^{2}}{2m_{c,y}}-\frac{V_{b}}{2}-\mu\right)$ and $\xi_{v,\mathbf{k}}=-\left(\frac{k_{x}^{2}}{2m_{v,x}}+\frac{k_{y}^{2}}{2m_{v,y}}-\frac{V_{b}}{2}+\mu\right)$, where $m_{c/v,x/y}$ are the anisotropic effective masses of the electrons in the conduction (top) and valence (bottom) layers, and $\mu$ is the chemical potential.
@@
Within each layer there is in-principle a two-fold spin degeneracy (or alternatively valley degeneracy as the valley is locked to spin via spin-orbit coupling [40]), which introduces additional possibilities for the spin-correlations of the exciton condensate. We circumvent this complication, by drawing motivation from a recent experimental study [42], where spin-polarized excitons have been reported in Coulomb-coupled monolayers of MoSe2 and WSe2; indeed mean-field studies of charge-compensated $\mathbf{Q}=0$} exciton condensates also indicate that spin-polarization is stabilized by an infinitesimal Zeeman field [74]. As such, we focus on an analogous situation, where a spin-polarized exciton condensate forms (for example, spin-$\uparrow$ from the conduction/top layer and spin-$\downarrow$ from the valence/bottom layer). In this sense, the spin-layer labels are ‘locked’. With this understanding, we henceforth drop the spin-labels and keep the layer labels in what follows. Although the spin-polarized exciton condensate breaks the underlying $C_{2z}$ symmetry, a combined operation ($\equiv C^{\prime}_{2z}$) of $C_{2z}$ and a $\pi$ spin-rotation about $\hat{x}$ remains a symmetry.
@@
$+\Big.2e^{-qd}n_{c}(\mathbf{q})n_{v}(-\mathbf{q})\Big],$ (1)
where $A$ is the area of the system, $V(\mathbf{q})=\frac{e^{2}}{2\epsilon_{0}\epsilon q}\text{tanh}(qd_{g})$ is the screened Coulomb interaction for a dual-gated setup with an inter-gate distance $d_{g}$, and interlayer separation $d$. We have compactly defined the conduction and valence densities, $n_{c/v}(\mathbf{q})=\sum_{\mathbf{k}}f^{\dagger}_{c/v,\mathbf{k}+\mathbf{q}}f_{c/v,\mathbf{k}}$. The system has a combined $U(1)_{c}\times U(1)_{v}$ symmetry, respectively corresponding to the independent conservation of particle number in each layer. The low-energy continuum model also possesses continuous translational symmetry, which reduces to discrete translational symmetries once a lattice potential is introduced. The interaction terms on the first line of Eq. Unconventional superconductivity mediated by exciton density wave fluctuations describe the intra-layer interactions, while the last term denotes the inter-layer interactions. Recent experiments [72, 39, 59, 42, 43] appear to realize inter-layer excitonic instabilties, which suggests that intra-layer interactions play the benign role of renormalizing the dispersion (and effective masses) and the native bandgap of the semiconductors; as such its role in triggering additional (within each layer) electronic instabilities is disregarded.
@@
Exciton density wave order.—The non-intersecting electron and hole Fermi surfaces in the respective top and bottom layers are susceptible to an exciton-formation instability of a finite-momentum (incommensurate) exciton condensation, as has been noted in previous works [50, 68]; the exciton condensed phase will be referred to as an exciton density wave (X-DW) [7]. In particular, with an unequal number of electron and hole carriers, the respective circular (elliptical) electron(hole) Fermi surfaces are separated by finite-momentum, which leads to the formation of finite-momentum excitons with a preferred center-of-mass momentum, $\mathbf{Q}$, as depicted in Fig. 1. The directional dependence of the center-of-mass momentum is explicitly confirmed [1], within a random phase approximation (RPA). The RPA susceptibility is peaked when the exciton center-of-mass momentum is $\pm\mathbf{Q}$, where $\mathbf{Q}=(k_{F,v,x}-k_{F,c,x})\hat{x}$ where $k_{F,v/c,x}$ are Fermi momenta in the two layers along the $x$-direction. With continuous rotational symmetry, exciton modes soften on a circle of points in momentum; the mass anisotropy lifts this degeneracy, leading to softening at discrete momenta $\pm\mathbf{Q}$.
Figure 2(a) depicts the Hartree-Fock phase diagram, which indicates the onset of exciton finite-momentum exciton instabilities, $\langle f^{\dagger}_{v,\mathbf{k}\pm\mathbf{Q}}f_{c,\mathbf{k}}\rangle$. In real-space, the exciton density wave order parameter is given by $\chi_{\mathbf{r}}(\mathbf{R})=\chi_{+}(\mathbf{R})e^{i\mathbf{Q}\cdot\mathbf{r}}+\chi_{-}(\mathbf{R})e^{-i\mathbf{Q}\cdot\mathbf{r}}$, where $\mathbf{r}$ is a center-of-mass coordinate and $\mathbf{R}$ is a relative coordinate between the electron and hole. As shown in Fig. 2(b), we find $\chi_{+}\approx\chi_{-}\equiv\chi_{0}$, so that the strength of the order parameter acquires a periodic spatial modulation in $\mathbf{r}$, with the phase staying uniform, resulting in the spontaneous breaking of $U(1)_{r}$ and continuous translational symmetry along the $x$-direction. We refer to this as the collinear X-DW. We note that the density range over which this collinear X-DW exists is for $|n|\lesssim 0.6\times 10^{12}$ cm-2 as shown in Fig. 2. We present in Fig. 3 a representative electronic bandstructure and Fermi surface depicting the impact of the exciton condensate, within a three-band Hubbard approximation (see Fig. A.1 in [1] for the self-consistent Hartree-Fock calculation that incorporates dual-gated Coulomb interactions). Importantly, despite a gap opening at certain momentum points in the Brillouin zone, a Fermi surface still remains.
@@
Figure 3: Development of collinear exciton density wave for $n=-0.7\times 10^{12}$ cm-2 and $V_{b}=12$ meV within a three-band Hubbard approximation (valence band for $\mathbf{k}\pm\mathbf{Q}$} and conduction band for $\mathbf{k}$}). Left: Electronic Fermi surface Fermi surface in the collinear X-DW phase (yellow denotes the filled Fermi sea). Right: Electronic Bandstructure for $k_{x}=0$ in the presence/absence of the collinear X-DW in red/black.
@@
Goldstone fluctuations mediated nodal superconductivity.—In the collinear X-DW condensate, breaking of the two continuous symmetries leads to two independent Goldstone modes. associated with: a uniform $U(1)_{r}$ phase rotation, $\chi=\chi_{0}\cos(\mathbf{Q}\cdot\mathbf{r})\rightarrow e^{i\theta}\chi_{0}\cos(\mathbf{Q}\cdot\mathbf{r})$, and a uniform translation $\chi=\chi_{0}\cos(\mathbf{Q}\cdot\mathbf{r})\rightarrow\chi_{0}\cos(\mathbf{Q}\cdot\mathbf{r}+\alpha)$. While the coupling of these Goldstone modes to the fermions retains the integrity of the Fermi liquid quasiparticles [73], we ask here whether they can mediate superconductivity. A similar theoretical study was performed in the context of inter-valley coherent fluctuation-mediated superconductivity in twisted bilayer graphene [51, 30] and multilayer graphene [14, 69]. For related works, see also Refs. [8, 26]; as well, Refs. [78, 70] for analyses of superconductivity arising from exciton fluctuations.
@@
We confine ourselves to the consideration of static fluctuations of the collinear X-DW order parameter: $\chi(\mathbf{r})=\chi_{0}\cos({\mathbf{Q}\cdot\mathbf{r}})+\delta\chi_{-}(\mathbf{r})e^{-i\mathbf{Q}\cdot\mathbf{r}}+\delta\chi_{+}(\mathbf{r})e^{i\mathbf{Q}\cdot\mathbf{r}}$, where the first term is the mean-field (saddle-point) solution. Retaining the important phase fluctuations ($\text{Im}[\delta\chi_{\pm}(\mathbf{r})]\neq 0$) at low energies, we obtain a Gaussian action for these fluctuations following the standard recipe of integrating out high-energy fermions (see [1] for details). Diagonalizing the $2\times 2$ fluctuation Hamiltonian, we obtain two gapless Goldstone modes: (i) ‘in-phase’ fluctuation mode ($\delta\chi_{+}=\delta\chi_{-}$, labelled by $\delta\chi_{s}$), which can be interpreted as a superfluid mode due to the breaking of $U(1)_{r}$, and (ii) an ‘out-of-phase’ mode ($\delta\chi_{+}=-\delta\chi_{-}$, labelled by $\delta\chi_{p}$), which can be interpreted as an acoustic phonon due to breaking of the continuous-translation symmetry by the fluctuations. The corresponding Goldstone mode Hamiltonian is,
$H_{G}=\frac{\rho_{G}}{2}\sum_{\mathbf{q}}(q_{x}^{2}+\kappa_{G}q_{y}^{2})\delta\chi_{G}(-\mathbf{q})\delta\chi_{G}(\mathbf{q}),$ (2)
@@
where $G=s/p$ labels the two (normal) Goldstone modes, $\rho_{G}$ is the Goldstone mode stiffness and $\kappa_{G}$ is the anisotropy parameter, both of which are numerically calculated. We note that we have disregarded the dynamics of the Goldstone mode here, which when treated provides a more accurate estimates of the critical temperature (within the framework of Eliashberg theory). Since our focus here is on the qualitative behaviors of the superconductivity, we focus solely on the static components; similar studies of fluctuation mediated superconductivity also demonstrate the predictive power of examining the purely static component of the Goldstone action [51, 30, 69]. The interactions induced by the Goldstone modes between the low-energy fermions $f$ (formed from the mean-field band) in the BCS channel is given by,
$\sum_{G}\frac{1}{2A\rho_{G}}\sum_{\mathbf{k},\mathbf{l}}V_{BCS}^{G}(\mathbf{k},\mathbf{l})\ f^{\dagger}_{\mathbf{k}}f^{\dagger}_{-\mathbf{k}}f_{-\mathbf{l}}f_{\mathbf{l}}.$ (3)
We provide details in the End Matter of how the Goldstone mode low-energy fermion coupling generates this attraction. Crucially, the superfluid mode’s interaction potential is even-parity under $C^{\prime}_{2z}$ ($V_{BCS}^{S}(\mathbf{k},\mathbf{l})=V_{BCS}^{S}(\mathbf{k},-\mathbf{l})$) and repulsive while the phonon mode has an odd-parity component ($V_{BCS}^{P}(\mathbf{k},\mathbf{l})\neq V_{BCS}^{P}(\mathbf{k},-\mathbf{l})$) and is attractive. As such, the Fermi-Dirac statistics enforced odd-parity superconducting parameter (of single-flavor fermions) can only be realized via the phonon mode.
@@
To analyze this possibility in detail, we self-consistently solve for the superconducting gap function using the interactions mediated by the phonon mode. Figure 4 depicts the superconducting order parameter (top) and the BCS gap function $\Delta(\mathbf{k})$ (bottom) of a representative parameter point obtained. The gap function has an anisotropic $p$-wave character, with nodal points in the BdG spectrum, and inherits the time-reversal symmetry breaking of the normal X-DW state. While time-reversal symmetry broken superconductors are uncommon in nature, there is recent evidence for this phenomenon in rhombohedrally-stacked tetra-layer graphene [22].
@@
Figure 4: Goldstone mode ($\delta\chi_{p}$)-mediated superconductivity inside the collinear X-DW phase. Top: Superconducting order parameter as a function of $V_{b}$ for $n=-0.7\times 10^{12}$ cm-2. Bottom: BCS gap function as a function of momentum for $V_{b}=12$ meV. Momentum mesh of $59\times 59$, $\epsilon=10$, and momentum space cutoff along the $x$-direction $k_{c,x}=0.2$.
@@
Critical fluctuations mediated pair-density wave.—The (mean-field suggested continuous) critical line demarcating the metallic electron-hole plasma and the X-DW in Fig. 2 indicates the development of further electronic instabilities. In the standard lore of metallic quantum critical points, superconductivity and non-Fermi liquids are two of the competing instabilities [36, 54]. Since our goal is not to study this competition in detail [41], we focus here on the possibility of unconventional superconductivity arising from pairing of the fermions on approach from the Fermi liquid. In [1], we consider non-Fermi liquid behavior from the classic one-loop hot-spot model treatment which leads to a dynamical critical exponent $z=3$, and singular corrections of the electronic self-energy, $\Sigma(\omega)\sim i|\omega|^{2/3}$.